Parth (parthkohli):

\[\begin{aligned}n&:=2 + 4 + 8 + 16 \cdots\\ \iff n&= 2(1 + 2 + 4 + 8 + 16)\cdots \\ \iff n &= 2(n + 1) \\ \iff n & = -1\end{aligned}\]

OpenStudy (anonymous):

the sum is a/r-1 2/2-1 2

OpenStudy (anonymous):

its an infinte gp series

OpenStudy (anonymous):

^yeah. and that's why you cannot subtract 2n and n to get n. It's like saying \[2\infty -\infty = \infty \] you just can't do that.